Evaluate.Round your answer to 3 decimal places.
what does the the answer to log (base 4, 12) mean?

Answer by stanbon(57407)

what does the the answer to log (base 4, 12) mean?
----
A "log" is an exponent or a power.
-----
Your problem is asking you to find the power of 4 needed to
get the number 12.
---
To get that answer you have to use a form called
"the change of base formula".
---
Here it is log4(12) = log(12)/log(4) = 1.0792/0.6021 = 1.7925
==========================
That means 4^1.7925 is very close to being 12.
Very close because we limited those numbers to 4 decimal places.
=======================
Cheers,
Stan H.